# Modeling Diffusion and Permeation Across the Stratum Corneum Lipid Barrier

### v1

@@ 1 Introduction
In 2024, the mean global near-surface temperature exceeded the pre-industrial average by more than $1.5\text{\,}\mathrm{K}$,1 underscoring the increasingly tangible impacts of global warming on human health, behavior, and daily life.2, 3, 4 A plausible behavioral response—particularly among vulnerable populations such as older adults and individuals with pre-existing health conditions—is to spend more time indoors to mitigate thermal discomfort during periods of extreme heat. However, empirical evidence remains limited regarding how climate change is reshaping patterns of time use.5, 6
@@
Nevertheless, because humans spend most of their time indoors, the impacts of climate change are largely mediated through indoor environments.7, 8 Accordingly, understanding how rising outdoor temperatures influence indoor air quality has become an increasingly important research priority.9, 10 Models and simulations are expected to play a central role in evaluating the long-term effects of climate change on indoor environments.11, 12 However, these modeling efforts rely on accurate physicochemical representations of the complex surface–air chemistry occurring indoors.13
@@
In this study, we address this challenge by investigating the permeation of two representative volatile and semivolatile organic compounds (VOCs) through human skin. The compounds of interest—acetone and 6-MHO (Figure 1)—are formed during the ozonolysis of skin lipids, which act as the dominant ozone scavengers in densely occupied indoor environments and thus constitute an important source of secondary indoor pollutants.14, 15 Kinetic models have been developed to describe the mass transport and chemical transformations of these oxidation products across the skin, surrounding air, and clothing.16, 17 However, the reliability of such models depends critically on accurate physicochemical parameters describing the partitioning of semivolatile compounds—such as 6-MHO—between the human body and indoor air,18 as well as on reliable descriptions of their diffusional transport through the skin barrier.
@@
To address this challenge, we recently began investigating permeation energetics using atomistic molecular dynamics (MD) simulations of stratum corneum (SC) lipid membranes (Fig 1)—the skin’s primary barrier to chemical penetration.19 We found that the convergence of free-energy (FE) profiles describing the translocation of permeants across these membranes was remarkably slow, requiring microseconds of simulation time even when state-of-the-art enhanced sampling techniques were employed. We traced the molecular origin of this convergence bottleneck to the highly ordered structure of SC membranes, which contrasts sharply with the fluid-like organization of conventional lipid bilayers such as those composed of 1-palmitoyl-2-oleoyl-sn-glycero-3-phosphocholine (POPC). In addition, we identified stochastic lipid flip–flop events between the two membrane leaflets that induce long-lived structural asymmetries.
@@
Figure 1: Chemical structures of acetone and 6-MHO, and a representative snapshot of the model system. The atomistic lipid membrane is composed of cholesterol (orange), lignoceric acid (yellow), and ceramide NS (green). In this study, we investigate solute permeation across the membrane along the $z$-axis. Figure adapted with permission from ref. 19.
@@
In this work, we extend our previous investigations by providing estimates of diffusivity profiles along the membrane normal. Predicting these diffusivities remains a challenging task—particularly when permeation is slow and enhanced sampling is required to overcome large FE barriers. We employ two complementary approaches to compute position-dependent diffusivities from harmonically restrained MD simulations, based on either the velocity autocorrelation function (VACF)20 or the position autocorrelation function (PACF)21 of the coordinate describing membrane permeation. As shown by our results—and consistent with the findings of Rowley et al. for a dipalmitoylphosphatidylcholine (DPPC) membrane22—estimates obtained from the two approaches differ substantially. Through a rigorous propagator-based analysis, we demonstrate that the resulting diffusivity profiles provide lower and upper bounds to the true profile. We further propagate these results to obtain single-membrane permeabilities and we critically assess how methodological uncertainties affect the accuracy of these predictions. Taken together, our work advances the understanding of molecular permeation across skin lipid membranes and establishes a robust framework for improving predictive models of indoor air chemistry under changing environmental conditions.
@@ 2 Methods
The model systems and simulation protocol have been described in detail in our previous study.19 Here, we briefly summarize the main aspects relevant to the present work and describe the computation of position-dependent diffusivities and the propagator analysis.
@@ 2.1 Model Systems
We studied three solutes—acetone, 6-MHO, and water—representing two skin-oil oxidation products and a reference compound. These solutes permeate through a model stratum corneum (SC) lipid bilayer, originally proposed by Klauda et al., and representative of the short periodicity phase (SPP)—the tightly packed lamellar arrangement of lipids that forms the primary diffusion barrier in the SC (Figure 1).23, 24
@@ 2.2 Umbrella Sampling
The translocation of the solutes across the membrane is described by a collective variable (CV), denoted as $z$, which measures the center-of-mass distance between each solute and the membrane projected onto the membrane normal (Figure 1).
For each solute, we performed 91 independent simulations (windows), in which the solute was harmonically restrained at integer-$\mathrm{\text{Å}}$ positions along the $z$ coordinate (umbrella sampling, US). A detailed analysis of the influence of the harmonic force constant was presented by Rowley et al.,22 and our own related findings on this topic are provided in Section S1 of the Supporting Information (SI). For the results reported in the main text, a force constant of $30\text{\,}\mathrm{\text{kcal}}\text{\,}{\mathrm{mol}}^{-1}\text{\,}{\mathrm{\text{Å}}}^{-2}$ is employed. Each simulation is initiated from well-equilibrated configurations obtained in our previous study and run for $15\text{\,}\mathrm{ns}$ per window.
[…]
@@ 2.3 Diffusivity Profiles
Two approaches are used to compute the position-dependent diffusivities, $D(z)$, along the membrane normal.
The first approach, proposed by Woolf and Roux, relies on the calculation of the velocity autocorrelation function (VACF):20
$C_{v}(t)=\left<\dot{z}(t)\dot{z}(0)\right$>, (1)
followed by the evaluation of a Laplace-frequency–dependent diffusivity,
@@
where $\hat{C}_{v}(s)$ denotes the Laplace transform of $C_{v}(t)$. The local diffusion coefficient is then obtained in the limit
$D(z=\left<z\right>)=\lim_{s\to 0}D(s),$ (3)
which, however, is mathematically ill-defined and therefore requires an empirical extrapolation procedure.
Hummer later proposed an alternative approach that avoids Laplace-domain extrapolation by using the position autocorrelation function (PACF):21
@@
In practice, however, one must either impose a cutoff time for the integral or assume a functional form to fit the correlation decay—both of which can introduce systematic bias.
The velocities required for computing the VACF are obtained from finite differences of the CV trajectory, which is saved at $2\text{\,}\mathrm{fs}$ intervals. Position-dependent diffusivities are determined using the ACFCalculator code developed by Rowley et al., which implements a heuristic procedure for the extrapolation in eq. 3.22
We note that alternative strategies for estimating position-dependent diffusivities are available. In particular, Bayesian-inference methods were originally proposed by Hummer28 and later extended to anisotropic media by Ghysels et al.29. They have also been coupled to adaptive biasing force calculations by Comer, Chipot, and co-workers.30, 31 Complementary formulations have been developed by Rosta and collaborators.32, 33 We focus here on the VACF/PACF approaches because they are widely used in membrane-permeation studies,34, 35, 36, 37, 38 straightforward to implement on the same umbrella-sampling data, and allow direct comparison to prior results for our benchmark system (water).
@@ 2.4 Propagator Analysis
where $\beta=1/(k_{\mathrm{B}}T)$ is the inverse thermal energy, $\Delta F(z)$ denotes the FE profile, and $D(z)$ represents the position-dependent diffusivity.
For given $\Delta F(z)$ and $D(z)$ profiles and a delta-function initial condition, $p(z,t_{0})=\delta(z-z_{0})$, this equation can be solved numerically using a standard forward-in-time, centered-in-space (FTCS) finite-difference scheme. The resulting solution, $p(z,t\,|\,z_{0},t_{0})$, represents the propagator of the diffusion model—that is, the conditional probability of finding a particle initially located at $z(t_{0})=z_{0}$ at position $z(t)$ at a later time $t$. The same quantity can also be obtained directly from molecular dynamics (MD) simulations without assuming diffusive behavior, thereby providing a rigorous test of the accuracy of the proposed $\Delta F(z)$ and $D(z)$ profiles.
@@ 3.1 Diffusivity Profiles
In the headgroup region, both methods indicate a local diffusivity bottleneck for all solutes. The most pronounced discrepancy between the two approaches, however, appears at the membrane center, where the estimates differ both quantitatively and qualitatively: the VACF method predicts higher-than-bulk diffusivities, whereas the PACF method yields lower values. Similar observations were reported by Rowley et al. for a DPPC membrane.22 In the following, we discuss the challenges associated with determining diffusivities using both approaches and the likely origins of these discrepancies. Our analysis focuses primarily on two representative regions of the membrane: the aqueous region, where the two methods show good agreement, and the membrane center, where the discrepancies are most pronounced.
@@ 3.2 Challenge 1: Extrapolation of the VACF Results
Figure 3: Illustration of determining diffusivities from velocity autocorrelation functions. The quantity $D(s)$, computed from eq. 2, must be extrapolated to zero by fitting a straight line over a range of $s$ values where the function remains well-behaved. This procedure performs reliably in the bulk phase but is highly sensitive to the placement of the tangent line for windows at the membrane center. The values obtained using the automated extrapolation procedure of Rowley et al.22 are indicated by the black and red points in Figure 2.
Obtaining position-dependent diffusivities from VACFs requires computing a Laplace-frequency–dependent diffusivity, $D(s)$, for each US window (eq. 2), followed by extrapolation to zero frequency (eq. 3). However, this limit is ill-defined because of the $s^{-1}$ dependence of $D(s)$ and therefore necessitates an empirical extrapolation procedure. Our VACF-based diffusivity profiles were obtained using the automated linear extrapolation method of Rowley et al.22 Although this approach is relatively robust in the aqueous bulk region, it may not be optimal at the membrane center (Figure 3). Because $D(s)$ exhibits pronounced curvature in the membrane core, the linear extrapolation is highly sensitive to the placement of the tangent line, yielding diffusivity estimates between roughly $0.1\text{\,}\mathrm{\text{Å}}^{2}\mathrm{/}\mathrm{p}\mathrm{s}$ and $0.75\text{\,}\mathrm{\text{Å}}^{2}\mathrm{/}\mathrm{p}\mathrm{s}$.

### v2

@@ 1 Introduction
In 2024, the mean global near-surface temperature exceeded the pre-industrial average by more than $1.5\text{\,}\mathrm{K}$,WorldMeteorologicalOrganizationWMO.WorldMeteorologicalOrganizationWMO.2025.StateGlobalClimate underscoring the increasingly tangible impacts of global warming on human health, behavior, and daily life.Rocque.Witteman.2021.HealthEffectsClimate, IntergovernmentalPanelOnClimateChangeIPCC.IntergovernmentalPanelOnClimateChangeIPCC.2022.Impacts15degCGlobal, Abdul-Nabi.Zahran.2025.ClimateChangeIts A plausible behavioral response—particularly among vulnerable populations such as older adults and individuals with pre-existing health conditions—is to spend more time indoors to mitigate thermal discomfort during periods of extreme heat. However, empirical evidence remains limited regarding how climate change is reshaping patterns of time use.White-Newsome.ONeill.2012.ClimateChangeHealth, Connolly.Connolly.2018.ClimateChangeAllocation
@@
Nevertheless, because humans spend most of their time indoors, the impacts of climate change are largely mediated through indoor environments.InstituteOfMedicine.InstituteOfMedicine.2011.ClimateChangeIndoor, Spengler.Spengler.2012.ClimateChangeIndoor Accordingly, understanding how rising outdoor temperatures influence indoor air quality has become an increasingly important research priority.Nazaroff.Nazaroff.2013.ExploringConsequencesClimate, Mansouri.Blondeau.2022.ImpactClimateChange Models and simulations are expected to play a central role in evaluating the long-term effects of climate change on indoor environments.Zhao.Schieweck.2024.LongtermPredictionEffects, Zhao.Hussein.2025.LongtermPredictionClimate However, these modeling efforts rely on accurate physicochemical representations of the complex surface–air chemistry occurring indoors.VonDomaros.Tobias.2025.MolecularDynamicsSimulations
@@
In this study, we address this challenge by investigating the permeation of two representative volatile and semivolatile organic compounds (VOCs) through human skin. The compounds of interest—acetone and 6-MHO (Figure 1)—are formed during the ozonolysis of skin lipids, which act as the dominant ozone scavengers in densely occupied indoor environments and thus constitute an important source of secondary indoor pollutants.Wisthaler.Weschler.2010.ReactionsOzoneHuman, Weschler.Weschler.2016.RolesHumanOccupant Kinetic models have been developed to describe the mass transport and chemical transformations of these oxidation products across the skin, surrounding air, and clothing.Lakey.Shiraiwa.2017.ChemicalKineticsMultiphase, Lakey.Shiraiwa.2019.ImpactClothingOzone However, the reliability of such models depends critically on accurate physicochemical parameters describing the partitioning of semivolatile compounds—such as 6-MHO—between the human body and indoor air,VonDomaros.Tobias.2020.MultiscaleModelingHuman as well as on reliable descriptions of their diffusional transport through the skin barrier.
@@
To address this challenge, we recently began investigating permeation energetics using atomistic molecular dynamics (MD) simulations of stratum corneum (SC) lipid membranes (Fig 1)—the skin’s primary barrier to chemical penetration.thomas_insights_2025 We found that the convergence of free-energy (FE) profiles describing the translocation of permeants across these membranes was remarkably slow, requiring microseconds of simulation time even when state-of-the-art enhanced sampling techniques were employed. We traced the molecular origin of this convergence bottleneck to the highly ordered structure of SC membranes, which contrasts sharply with the fluid-like organization of conventional lipid bilayers such as those composed of 1-palmitoyl-2-oleoyl-sn-glycero-3-phosphocholine (POPC). In addition, we identified stochastic lipid flip–flop events between the two membrane leaflets that induce long-lived structural asymmetries.
@@
Figure 1: Chemical structures of acetone and 6-MHO, and a representative snapshot of the model system. The atomistic lipid membrane is composed of cholesterol (orange), lignoceric acid (yellow), and ceramide NS (green). In this study, we investigate solute permeation across the membrane along the $z$-axis. Figure adapted with permission from ref. thomas_insights_2025.
@@
In this work, we extend our previous investigations by providing estimates of diffusivity profiles along the membrane normal. Predicting these diffusivities remains a challenging task—particularly when permeation is slow and enhanced sampling is required to overcome large FE barriers. We employ two complementary approaches to compute position-dependent diffusivities from harmonically restrained MD simulations, based on either the velocity autocorrelation function (VACF)Woolf.Roux.1994.ConformationalFlexibilityOPhosphorylcholine or the position autocorrelation function (PACF)Hummer.Hummer.2005.PositiondependentDiffusionCoefficients of the coordinate describing membrane permeation. As shown by our results—and consistent with the findings of Rowley et al. for a dipalmitoylphosphatidylcholine (DPPC) membranegaalswyk_generalized_2016—estimates obtained from the two approaches differ substantially. Through a rigorous propagator-based analysis, we demonstrate that the resulting diffusivity profiles provide lower and upper bounds to the true profile. We further propagate these results to obtain single-membrane permeabilities and we critically assess how methodological uncertainties affect the accuracy of these predictions. Taken together, our work advances the understanding of molecular permeation across skin lipid membranes and establishes a robust framework for improving predictive models of indoor air chemistry under changing environmental conditions.
@@ 2 Methods
We use atomistic MD simulations, which follow the motion of all atoms in the system in time according to classical mechanics. This approach allows us to quantify how individual molecules move through a model skin–lipid barrier and how readily they cross it. The resulting microscopic transport properties serve as input parameters for larger-scale models of skin transport and indoor air chemistry. The model systems and simulation protocol have been described in detail in our previous study.thomas_insights_2025 Here, we briefly summarize the main aspects relevant to the present work and describe the computation of position-dependent diffusivities and the propagator analysis.
@@ 2.1 Model Systems
The SC is a complex structure composed of cornified cells embedded in a lipid matrix, an arrangement often described as a brick-and-mortar architecture.Michaels.Shaw.1975.DrugPermeationHuman, Elias.Elias.1983.EpidermalLipidsBarrier, Schmitt.Neubert.2020.StateArtStratum Atomistic studies of skin permeation typically focus on the lipid matrix, which is the only continuous component of the SC and represents a major barrier to permeation. Extrapolation to the full SC requires combining MD-based transport properties with macroscopic models that account for additional SC components and the detailed permeation pathway.Wang.Nitsche.2007.MultiphaseMicroscopicDiffusion, Wang.Nagel.2023.TransdermalLateralEffective
The structure of the SC lipid matrix has been extensively characterized using a range of experimental techniques. There is broad consensus that it consists predominantly of an approximately equimolar mixture of ceramides (CERs), free fatty acids (FFAs), and cholesterol (CHOL), organized into at least two lamellar microphases: the short periodicity phase (SPP) and the long periodicity phase (LPP).Sparr.Topgaard.2023.StratumCorneumBarrier, Bouwstra.Gooris.2023.SkinBarrierExtraordinary
[…]
Here, we investigate the permeation of three solutes—acetone, 6-MHO, and water—representing two skin-oil oxidation products and a reference molecule. These solutes permeate through a model SC lipid bilayer originally proposed by Wang and Klaudawang_models_2018 and representative of the SPP. The present work focuses on the SPP as a tractable and widely used model system. Establishing robust methodology for diffusivity estimation in this simpler context provides a foundation for future studies that will incorporate the LPP and more realistic SC architectures.
@@ 2.2 Umbrella Sampling
The translocation of the solutes across the membrane is described by a collective variable (CV), denoted as $z$, which measures the center-of-mass distance between each solute and the membrane projected onto the membrane normal (Figure 1). The CV $z$ can be viewed as a one-dimensional reaction coordinate that tracks how far the molecule has moved from the membrane center toward either side. Instead of relying on a single long simulation to capture rare permeation events, we divide the coordinate into many overlapping windows along $z$, where each window focuses the sampling on a narrow region of the membrane, in a procedure known as umbrella sampling (US). The solute is confined within each window by a harmonic potential, which permits fluctuations around a chosen position while still allowing the solute to explore the local environment. For each solute, we performed 91 US windows, in which the solute was harmonically restrained at integer-$\mathrm{\text{Å}}$ positions along the $z$ coordinate.
A detailed analysis of the influence of the harmonic force constant was presented by Rowley et al.,gaalswyk_generalized_2016 and our own related findings on this topic are provided in Section S1 of the Supporting Information (SI). For the results reported in the main text, a force constant of $30\text{\,}\mathrm{\text{kcal}}\text{\,}{\mathrm{mol}}^{-1}\text{\,}{\mathrm{\text{Å}}}^{-2}$ is employed. Each simulation is initiated from well-equilibrated configurations obtained in our previous study and run for $15\text{\,}\mathrm{ns}$ per window. Although simulation times per window are short, the combined sampling over all windows effectively corresponds to many independent permeation attempts across the entire membrane. All simulations are carried out at a temperature close to that of the outer skin surface ($T=$305.15\text{\,}\mathrm{K}$\$) to mimic realistic physiological conditions. The temperature is maintained by a stochastic velocity-rescaling thermostatbussi_canonical_2007 with a time constant of $1\text{\,}\mathrm{ps}$. All simulations are performed using NAMD 3.0 (alpha13).phillips_scalable_2005, phillips_scalable_2020
@@ 2.3 Diffusivity Profiles
In addition to the FE barriers, a key quantity for transport is how quickly a molecule can move at different depths within the membrane. We express this by a position-dependent diffusivity $D(z)$, which may be interpreted as a local mobility: large $D(z)$ indicates that the molecule can move easily at that position, whereas small $D(z)$ reflects strong hindrance by the surrounding lipids. Two approaches are used to compute the position-dependent diffusivities, $D(z)$, along the membrane normal.
The first approach, proposed by Woolf and Roux, relies on the calculation of the velocity autocorrelation function (VACF):Woolf.Roux.1994.ConformationalFlexibilityOPhosphorylcholine
$C_{v}(t)=\left<\dot{z}(t)\dot{z}(0)\right$>. (1)
The VACF quantifies how quickly the solute’s velocity loses memory of its initial direction of motion. If a molecule continues to move in approximately the same direction for a long time, it is effectively more mobile than if its velocity becomes randomized on a short timescale. To extract a position-dependent diffusivity $D(z)$, one must perform a Laplace transform of the VACF, $\hat{C}_{v}(s)$, and extrapolate the resulting Laplace-frequency–dependent diffusivity,
@@
to the limit of zero frequency (infinite lag time),
$D(z=\left<z\right>)=\lim_{s\to 0}D(s).$ (3)
However, this limit is ill-defined because of the $s^{-1}$ dependence and therefore requires an empirical extrapolation procedure.
Hummer later proposed an alternative approach that avoids Laplace-domain extrapolation by using the position autocorrelation function (PACF):Hummer.Hummer.2005.PositiondependentDiffusionCoefficients
@@
The PACF characterizes how strongly the current position of the solute along $z$ remains correlated with its initial position. In regions where the molecule moves freely, this positional correlation decays rapidly, whereas in regions where its motion is constrained, the correlation persists for longer times. In practice, however, one must either impose a cutoff time for the integral or assume a functional form to fit the correlation decay—both of which can introduce systematic bias.
The velocities required for computing the VACF are obtained from finite differences of the CV trajectory, which is saved at $2\text{\,}\mathrm{fs}$ intervals. Position-dependent diffusivities are determined using the ACFCalculator code developed by Rowley et al., which implements a heuristic procedure for the extrapolation in eq. 3.gaalswyk_generalized_2016
We note that alternative strategies for estimating position-dependent diffusivities are available. In particular, Bayesian-inference methods, which were originally proposed by Hummerhummer_position-dependent_2005 and later extended to anisotropic media by Ghysels et al.Ghysels.Hummer.2017.PositionDependentDiffusionTensors, have also been developed. They have also been coupled to adaptive biasing force calculations by Comer, Chipot, and co-workers.Comer.Gonzalez-Nilo.2013.CalculatingPositionDependentDiffusivity, Chipot.Comer.2016.SubdiffusionMembranePermeation Complementary formulations have been developed by Rosta and collaborators.Rosta.Hummer.2015.FreeEnergiesDynamic, Sicard.Rosta.2021.PositionDependentDiffusionBiased Position-dependent diffusivities can also be estimated from the autocorrelation function of the constraint force (FACF) in simulations where the particle is algorithmically constrained to selected positions $z$ along the CV. However, this approach can be sensitive to finite-size and momentum-conservation artifacts and therefore requires particular care in its practical implementation.FujimotoNagaiEtAl2021 We focus here on the VACF/PACF approaches because they are widely used in membrane permeation studies,das_water_2009, gupta_molecular_2016, kramer_membrane_2020, piasentin_evaluation_2021, Reuter.Lunter.2025.PresenceDifferentCeramide and they allow direct comparison with prior results for our benchmark system (water).
Using both VACF- and PACF-based estimators on the same simulation data allows us to assess how sensitive the inferred diffusivities are to the choice of analysis method. As shown below, the two approaches agree well in the bulk water regions but diverge near the membrane center, where the dynamics is more complex. We interpret these differences as providing a realistic range (lower and upper bounds) for the true local diffusivity.
@@ 2.4 Propagator Analysis
where $\beta=1/(k_{\mathrm{B}}T)$ is the inverse thermal energy, $\Delta F(z)$ denotes the FE profile, and $D(z)$ represents the position-dependent diffusivity. This one-dimensional Smoluchowski equation describes diffusion on a free-energy landscape: the molecule tends to move from regions of high free energy to regions of lower free energy, while its motion is locally slowed down or accelerated by the position-dependent diffusivity $D(z)$. In this formulation, the thermodynamic driving forces are encoded in $\Delta F(z)$ and the kinetic hindrance in $D(z)$, and both contributions enter a single evolution equation for the probability of finding the molecule at a given position and time.
For given $\Delta F(z)$ and $D(z)$ profiles and a delta-function initial condition, $p(z,t_{0})=\delta(z-z_{0})$, this equation can be solved numerically using a standard forward-in-time, centered-in-space (FTCS) finite-difference scheme. The resulting solution, $p(z,t\,|\,z_{0},t_{0})$, represents the propagator of the diffusion model—that is, the conditional probability of finding a particle initially located at $z(t_{0})=z_{0}$ at position $z(t)$ at a later time $t$. The same quantity can also be obtained directly from molecular dynamics (MD) simulations without assuming diffusive behavior, thereby providing a rigorous test of the accuracy of the proposed $\Delta F(z)$ and $D(z)$ profiles. Comparing these model-based propagators to those obtained directly from MD simulations provides a stringent test of the validity of the chosen $\Delta F(z)$ and $D(z)$ profiles.
@@ 3.1 Diffusivity Profiles
In the headgroup region, both methods indicate a local diffusivity bottleneck for all solutes. The most pronounced discrepancy between the two approaches, however, appears at the membrane center, where the estimates differ both quantitatively and qualitatively: the VACF method predicts higher-than-bulk diffusivities, whereas the PACF method yields lower values. Similar observations were reported by Rowley et al. for a DPPC membrane.gaalswyk_generalized_2016 In the following, we discuss the challenges associated with determining diffusivities using both approaches and the likely origins of these discrepancies. Our analysis focuses primarily on two representative regions of the membrane: the aqueous region, where the two methods show good agreement, and the membrane center, where the discrepancies are most pronounced.
@@ 3.2 Challenge 1: Extrapolation of the VACF Results
Figure 3: Illustration of determining diffusivities from velocity autocorrelation functions. The quantity $D(s)$, computed from eq. 2, must be extrapolated to zero by fitting a straight line over a range of $s$ values where the function remains well-behaved. This procedure performs reliably in the bulk phase but is highly sensitive to the placement of the tangent line for windows at the membrane center. The values obtained using the automated extrapolation procedure of Rowley et al.gaalswyk_generalized_2016 are indicated by the black and red points in Figure 2.
Obtaining position-dependent diffusivities from VACFs requires computing a Laplace-frequency–dependent diffusivity, $D(s)$, for each US window (eq. 2), followed by extrapolation to zero frequency (eq. 3). However, this limit is ill-defined because of the $s^{-1}$ dependence of $D(s)$ and therefore necessitates an empirical extrapolation procedure. Our VACF-based diffusivity profiles were obtained using the automated linear extrapolation method of Rowley et al.gaalswyk_generalized_2016 Although this approach is relatively robust in the aqueous bulk region, it may not be optimal at the membrane center (Figure 3). Because $D(s)$ exhibits pronounced curvature in the membrane core, the linear extrapolation is highly sensitive to the placement of the tangent line, yielding diffusivity estimates between roughly $0.1\text{\,}\mathrm{\text{Å}}^{2}\mathrm{/}\mathrm{p}\mathrm{s}$ and $0.75\text{\,}\mathrm{\text{Å}}^{2}\mathrm{/}\mathrm{p}\mathrm{s}$.